# VideoGen-of-Thought: Step-by-step generating multi-shot video with minimal manual intervention

### v2

@@ Abstract
Current video generation models excel at short clips but fail to produce cohesive multi-shot narratives due to disjointed visual dynamics and fractured storylines. Existing solutions either rely on extensive manual scripting/editing or prioritize single-shot fidelity over cross-scene continuity, limiting their practicality for movie-like content. We introduce VideoGen-of-Thought (VGoT), a step-by-step framework that automates multi-shot video synthesis from a single sentence by systematically addressing three core challenges: (1) Narrative Fragmentation: Existing methods lack structured storytelling. We propose dynamic storyline modeling, which first converts the user prompt into concise shot descriptions, then elaborates them into detailed, cinematic specifications across five domains (character dynamics, background continuity, relationship evolution, camera movements, HDR lighting), ensuring logical narrative progression with self-validation. (2) Visual Inconsistency: Existing approaches struggle with maintaining visual consistency across shots. Our identity-aware cross-shot propagation generates identity-preserving portrait (IPP) tokens that maintain character fidelity while allowing trait variations (expressions, aging) dictated by the storyline. (3) Transition Artifacts: Abrupt shot changes disrupt immersion. Our adjacent latent transition mechanisms implement boundary-aware reset strategies that process adjacent shots’ features at transition points, enabling seamless visual flow while preserving narrative continuity. By integrating these innovations into a training-free pipeline, VGoT generates multi-shot videos that outperform state-of-the-art baselines by 20.4% in within-shot face consistency and 17.4% in style consistency, while achieving over 100% better cross-shot consistency and 10× fewer manual adjustments than alternatives like MovieDreamer and DreamFactory. Our work redefines automated long-video generation, bridging the gap between raw visual synthesis and director-level storytelling.
@@ 1 Introduction
Recent advancements in video generation techniques have yielded impressive results, particularly in creating short, visually appealing clips [5, 7, 8, 53, 24]. These advancements have been powered by increasingly sophisticated generative models, ranging from diffusion models [25, 46, 40, 5] to auto-regressive models [15, 49, 33, 48], supported by large-scale datasets [28, 41, 42]. These methods have enabled the generation of high-quality and realistic short videos. However, generating multi-shot videos from a brief user input script remains a substantial challenge. Unlike single-shot video generation, which focuses on creating a coherent clip from a single prompt, multi-shot video generation requires the model to maintain both reasonable storylines and visual consistency across multiple shots. This task involves additional complexities, such as ensuring logical transitions between scenes and maintaining consistent appearance (e.g., character identity, overall style, etc.) throughout the video. Current video generation methods [7, 8, 53, 27, 54] often fall short in these areas, resulting in fragmented narratives and inconsistent visual elements across shots. Furthermore, sometimes real movies require the same character to appear in different ways based on the storylines, which should be faithful to the same identity but not the same traits (e.g., expression, appearance, relationships, etc.). The requirement for high-level identity preservation [55, 57, 62] across shots remains an open question, which is essential for cross-shot consistency.
Existing multi-shot video generation approaches suffer from several limitations. For instance, methods like MovieDreamer [60] require plenty of manual input, including mountains of script writing (e.g., character appearance, scene elements, detailed plots, etc.) and image selection. Other approaches, such as DreamFactory [51], focus on multi-agent pipelines but require specific documents and repeated manual adjustment for each story, restricting the capability of easy usage. The need for heavy manual intervention not only increases the workload but also limits their practicality for automated movie-like content creation. In contrast, our work aims to decompose the complex task of multi-shot video generation into smaller, manageable problems and solve them in a step-by-step manner with minimal manual intervention, as shown in Figure 1. Our approach proposes to systematically and automatically address three core challenges: (1) narrative fragmentation through dynamic storyline modeling; (2) visual inconsistency via identity-aware cross-shot propagation; and (3) transition artifacts using adjacent latent transition mechanisms.
@@
Additionally, to resolve visual inconsistency, we introduce identity-aware cross-shot propagation that extracts multi-aspect portrait schemata from narrative elements, ranging from different avatars in the same story to different traits of the same identity following the development of the story. This system handles both inter-story avatar variations and intra-story identity evolution, generating identity-preserving portrait (IPP) tokens. These IPP tokens guide keyframe generation through hierarchical feature injection in pretrained diffusion models [45, 40, 47, 55], ensuring style uniformity and identity fidelity across shots.
We encode each keyframe into latent space and utilize a video diffusion model to refine a noise map into video latent codes, representing $k$ frames of shot video, conditioned on the keyframe latent and the corresponding textual latent. To address transition artifacts, we design a cross-shot transition mechanism with a FIFO-like [30] latent reset strategy that processes adjacent shots’ features at the boundary, ensuring seamless transitions and maintaining visual continuity across shots while preserving the logical coherence established in the storyline preparation.
@@
Experimental results demonstrate VGoT’s superiority over state-of-the-art (SOTA) methods across all metrics. Quantitative comparisons reveal 20.4% and 17.4% improvements in WS-FC and WS-SC, respectively, compared to previous SOTA baselines. For cross-shot metrics, VGoT achieves 2.9$\times$ higher in CS-FC and 106.6% higher in CS-SC over baselines. Human evaluations (Table 2) confirm these findings, with VGoT receiving 66.7% ”Good” ratings for cross-shot consistency versus 27.2% for competitors.
@@
  • Three-Core Solution Architecture: We propose a structured four-module pipeline addressing
    1. 1.
      Narrative Fragmentation through LLM-powered story decomposition
    2. 2.
      Visual Inconsistency via identity-aware propagation
    3. 3.
      Transition Artifacts using adjacent latent transition mechanisms
@@ 2 Related Work
Long Video Synthesis. Long video synthesis has advanced through diffusion-based methods and autoregressive approaches. Diffusion models based on Stable Diffusion [40] utilize iterative refinement to generate visually consistent frames and have been effective for short sequences [22, 5, 7, 8, 52, 50, 6, 21, 54, 59, 17, 29, 36, 13, 26, 44, 58, 61, 37, 35]. However, they struggle with maintaining coherence over extended video lengths. Autoregressive methods [15, 32, 56] predict frames sequentially but often face error accumulation, making long-term consistency challenging and computationally expensive. Additionally, training-free methods like FIFO-Diffusion [30] generate long sequences without training but lack mechanisms to manage transitions across shots, limiting their effectiveness in narrative-driven content. Overall, while these approaches achieve visual fidelity, they fail to ensure logical coherence across extended sequences. In contrast, VideoGen-of-Thought (VGoT) leverages a modular approach that includes cross-shot smoothing mechanisms to ensure both visual consistency and narrative coherence, offering a more holistic solution for generating long-form videos.
Multi-Shot Video Generation. Multiple Shot Story Generation focuses on maintaining narrative coherency across distinct scenes, and existing approaches face critical limitations in automation scalability. Animate-a-Story [23] uses retrieval-augmented generation to ensure visual consistency but struggles with maintaining logical narrative transitions. MovieDreamer [60] requires extensive manual scripting (character profiles, scene details, etc.) and image curation. DreamFactory [51] demands repetitive adjustment and special documents for each story in its multi-agent system across stories. Flexifilm [34] and StoryDiffusion [62] introduce conditional adaptability but still necessitate manual consistency fixes between shots. These methods share a common bottleneck: heavy reliance on human intervention for narrative and visual coherence. VGoT breaks this paradigm through systematic automation: 1) Narrative Fragmentation → LLM-driven story decomposition; 2) Visual Inconsistency → Story-derived identity propagation; 3) Transition Artifacts → Boundary-aware latent processing. This structured approach enables director-level storytelling from single prompts with 10× fewer manual interventions than prior works [60, 51].
@@ 3.1 Preliminaries
Diffusion models [25, 45, 46] are generative models trained to approximate data distributions $p(x)$ through iterative denoising of random noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. The forward process gradually adds noise according to a variance schedule $\beta_{t}$:
@@
Latent diffusion models [40] map this process to a compressed space using a VAE [31], enabling conditional generation through cross-attention:
@@
where $Q=W_{Q}\mathcal{E}(x_{t})$, $K=W_{K}\tau_{\theta}(y)$, and $V=W_{V}\tau_{\theta}(y)$ for text embeddings $y$. Video diffusion models [22, 5] process frame sequences $\{z^{f}\}_{f=0}^{F-1}\in\mathbb{R}^{h\times w\times d}$ through temporal-aware denoising networks $\epsilon_{\theta}(z_{t},t,c)$. FIFO-Diffusion [30] extends this through queue-based latent processing:
@@
where $\Phi$ denotes the DDIM sampler [45]. While effective for frame continuity, this approach struggles with: (1) abrupt shot transitions that disrupt queue coherence, and (2) integration of image-based conditions. Our framework addresses these through reset mechanisms in adjacent latent-space processing.
@@ 3.2 Problem Definition
Given a one-sentence user input $S$ specifying $N$ shots (e.g., ”A story of Mary’s life from birth to death”), we aim to generate a multi-shot video $V$ with reasonable storylines and visual consistency from one sentence through minimal manual intervention. The core challenges are:
@@
  • Consistency: Preserving high-level identity [55, 57, 62] while allowing trait variations (e.g., aging expressions, contextual relationships) across shots
@@ 4 Method: VideoGen-of-Thought
Our methodology introduces a structured, step-by-step framework for generating multi-shot videos with reasonable storylines and visual consistency from one sentence with minimal manual intervention, addressing the core challenges of narrative fragmentation, visual inconsistency, and transition artifacts through four distinct yet collaborative modules (Fig 2):
@@ 4.1 Dynamic Storyline Modeling with Self-Vaildation
where $\mathcal{C}_{\text{film}}\triangleq(p_{\text{cha}},p_{b},p_{r},p_{\text{cam}},p_{h})$ encodes five cinematic dimensions: $p_{\text{cha}}$ governs character appearance evolution and role relationships, $p_{b}$ ensures background consistency across scene transitions, $p_{r}$ maintains interaction patterns and event causality, $p_{\text{cam}}$ specifies shot composition through camera movements, and $p_{h}$ regulates HDR lighting continuity. The self-validation mechanism employs two criteria:
@@
where $C:\mathcal{P}\times\mathcal{P}\rightarrow[0,1]$ measures narrative coherence through pretrained textual feature extractor $E_{\text{GLM}}$ [18] to compute semantic similarity between consecutive shots, and $K:\mathcal{P}\times\mathcal{C}\rightarrow\{0,1\}$ verifies constraint completeness via rule-based checks against $\mathcal{C}_{\text{film}}$. Thresholds $\tau_{c}=0.85$ and $\tau_{k}=1$ ensure strict adherence to cinematic principles.
@@ 4.2 Identity-Aware Cross-Shot Propagation
where $\mathcal{M}_{I}$ is a pre-trained text-to-image model and $E_{\text{CLIP}}$ denotes CLIP [38]’s vision encoder. We inject identity features into diffusion via cross-attention:
@@
where $z_{t}\in\mathbb{R}^{n\times d}$ is the latent in step $t$, $\lambda$ balances the influence of text / identity, and [;][;][ ; ] denotes concatenation. Keyframe generation integrates both modalities:
@@ 4.3 Adjacent Latent Transition Mechanisms
where $e^{T}_{i}=E_{\text{GLM}}(s_{i})$ uses simplified shot description $s_{i}$ rather than detailed script $p_{i}$, $e^{I}_{i}$ is the keyframe embedding, and $\epsilon_{i}\sim\mathcal{N}(0,\mathbf{I})$ is the initial noise. Inspired by FIFO [30], we implement boundary-aware noise reset for cross-shot transitions:
@@
EasyAnimate [53] 0.2402 0.4705 0.0268 0.7969 0.2037
CogVideo [54] 0.2477 0.6099 0.0222 0.7424 0.2069
VideoCrafter1 [7] 0.2478 0.3706 0.0350 0.7623 0.1867
VideoCrafter2 [8] 0.2529 0.5569 0.0686 0.7981 0.1798
@@
EasyAnimate [53] 0.3333 0.3232 0.3434 0.3535 0.3535 0.3131 0.4646 0.2727 0.2828
CogVideo [54] 0.1341 0.4146 0.4512 0.2927 0.5976 0.2317 0.1463 0.4512 0.5244
VideoCrafter1 [7] 0.5446 0.2574 0.1980 0.6436 0.1881 0.1683 0.6535 0.1782 0.1683
VideoCrafter2 [8] 0.1262 0.4854 0.3883 0.3495 0.3786 0.2718 0.1748 0.4951 0.3981
@@
Figure 4: Visual comparison of VGoT and baselines
4.4 Multi-Shot Evaluation Protocol
where $F_{j}^{i}$ represents facial features from frame $j$ of shot $i$ extracted using InsightFace [39, 20, 16, 3, 2, 11, 12, 19, 9, 10].
Within-Shot Style Consistency (WS-SC) quantifies stylistic coherence through VGG-19 [43] features:
where $S_{j}^{i}$ represents flattened VGG-19 [43] features from frame $j$ of shot $i$.

### v3

# What Makes a Good Video: Next Practices in Video Generation and Evaluation

@@ Abstract
Current video generation models excel at short clips but fail to produce cohesive multi-shot narratives due to disjointed visual dynamics and fractured storylines. Existing solutions either rely on extensive manual scripting/editing or prioritize single-shot fidelity over cross-scene continuity, limiting their practicality for movie-like content. We introduce VideoGen-of-Thought (VGoT), a step-by-step framework that automates multi-shot video synthesis from a single sentence by systematically addressing three core challenges: (1) Narrative fragmentation: Existing methods lack structured storytelling. We propose dynamic storyline modeling, which turns the user prompt into concise shot drafts and then expands them into detailed specifications across five domains (character dynamics, background continuity, relationship evolution, camera movements, and HDR lighting) with self-validation to ensure logical progress. (2) Visual inconsistency: previous approaches struggle to maintain consistent appearance across shots. Our identity-aware cross-shot propagation builds identity-preserving portrait (IPP) tokens that keep character identity while allowing controlled trait changes (expressions, aging) required by the story. (3) Transition artifacts: Abrupt shot changes disrupt immersion. Our adjacent latent transition mechanisms implement boundary-aware reset strategies that process adjacent shots’ features at transition points, enabling seamless visual flow while preserving narrative continuity. Combined in a training-free pipeline, VGoT surpasses strong baselines by 20.4% in within-shot face consistency and 17.4% in style consistency, while requiring 10× fewer manual adjustments. VGoT bridges the gap between raw visual synthesis and director-level storytelling for automated multi-shot video generation.
@@
footnotetext: † Corresponding author.footnotetext: Project webpage: https://cheliosoops.github.io/VGoT/
@@ 1 Introduction
Recent advancements in video generation techniques have yielded impressive results, particularly in creating short, visually appealing clips Blattmann et al. (2023a); Chen et al. (2023, 2024); Xu et al. (2024); Henschel et al. (2024). These advancements have been powered by increasingly sophisticated generative models, ranging from diffusion models Ho et al. (2020); Song et al. (2020b); Rombach et al. (2022); Blattmann et al. (2023a) to auto-regressive models Ge et al. (2022); Weng et al. (2023); Liu et al. (2024); Wang et al. (2024), supported by large-scale datasets Huang et al. (2020); Schuhmann et al. (2021, 2022). These methods have enabled the generation of high-quality and realistic short videos. However, generating multi-shot videos from a brief user input script remains a substantial challenge. Unlike single-shot video generation, which focuses on creating a coherent clip from a single prompt, multi-shot video generation requires the model to maintain both reasonable storylines and visual consistency across multiple shots. This task involves additional complexities, such as ensuring logical transitions between scenes and maintaining consistent appearance (e.g., character identity, overall style, etc.) throughout the video. Current video generation methods Chen et al. (2023, 2024); Xu et al. (2024); Hong et al. (2022); Yang et al. (2024) often fall short in these areas, resulting in fragmented narratives and inconsistent visual elements across shots. Furthermore, sometimes real movies require the same character to appear in different ways based on the storylines, which should be faithful to the same identity but not the same traits (e.g., expression, appearance, relationships, etc.). The requirement for high-level identity preservation Ye et al. (2023); Yuan et al. (2024); Zhou et al. (2025) across shots remains an open question, which is essential for cross-shot consistency.
Existing multi-shot video generation approaches suffer from several limitations. For instance, methods like MovieDreamer Zhao et al. (2024) require plenty of manual input, including mountains of script writing (e.g., character appearance, scene elements, detailed plots, etc.) and image selection. Other approaches, such as DreamFactory Xie et al. (2024), focus on multi-agent pipelines but require specific documents and repeated manual adjustment for each story, restricting the capability of easy usage. The need for heavy manual intervention not only increases the workload but also limits their practicality for automated movie-like content creation. In contrast, our work aims to decompose the complex task of multi-shot video generation into smaller, manageable problems and solve them in a step-by-step manner with minimal manual intervention, as shown in Figure 1. Our approach proposes to systematically and automatically address three core challenges: (1) narrative fragmentation through dynamic storyline modeling; (2) visual inconsistency via identity-aware cross-shot propagation; and (3) transition artifacts using adjacent latent transition mechanisms.
@@
Additionally, to resolve visual inconsistency, we introduce identity-aware cross-shot propagation that extracts multi-aspect portrait schemata from narrative elements, ranging from different avatars in the same story to different traits of the same identity following the development of the story. This system handles both inter-story avatar variations and intra-story identity evolution, generating identity-preserving portrait (IPP) tokens. These IPP tokens guide keyframe generation through hierarchical feature injection in pretrained diffusion models Song et al. (2020a); Rombach et al. (2022); Team (2024); Ye et al. (2023), ensuring style uniformity and identity fidelity across shots.
We encode each keyframe into latent space and utilize a video diffusion model to refine a noise map into video latent codes, representing $k$ frames of shot video, conditioned on the keyframe latent and the corresponding textual latent. To address transition artifacts, we design a cross-shot transition mechanism with a FIFO-like Kim et al. (2024) latent reset strategy that processes adjacent shots’ features at the boundary, ensuring seamless transitions and maintaining visual continuity across shots while preserving the logical coherence established in the storyline preparation.
@@
[…]
@@
  • Three-Core Solution Architecture: We propose a structured four-module pipeline addressing narrative fragmentation through LLM-powered story decomposition, visual inconsistency via identity-aware propagation, and transition artifacts using adjacent latent transition mechanisms.
@@ 2 Related Work
Long Video Synthesis. Long video synthesis has advanced through diffusion-based methods and autoregressive approaches. Diffusion models based on Stable Diffusion Rombach et al. (2022) utilize iterative refinement to generate visually consistent frames and have been effective for short sequences He et al. (2022); Blattmann et al. (2023a); Chen et al. (2023, 2024); Xing et al. (2025); Wu et al. (2023); Blattmann et al. (2023b); Guo et al. (2023); Yang et al. (2024); Zhang et al. (2024); Geyer et al. (2023); Huang et al. (2024); Peng et al. (2024); Esser et al. (2023); Ho et al. (2022); Singer et al. (2022); Zeng et al. (2024); Zhou et al. (2022); Qiu et al. (2023); Pan et al. (2024). However, they struggle with maintaining coherence over extended video lengths. Autoregressive methods Ge et al. (2022); Li et al. (2024); Yin et al. (2023) predict frames sequentially but often face error accumulation, making long-term consistency challenging and computationally expensive. Additionally, training-free methods like FIFO-Diffusion Kim et al. (2024) generate long sequences without training but lack mechanisms to manage transitions across shots, limiting their effectiveness in narrative-driven content. Overall, while these approaches achieve visual fidelity, they fail to ensure logical coherence across extended sequences. In contrast, VideoGen-of-Thought (VGoT) leverages a modular approach that includes cross-shot smoothing mechanisms to ensure both visual consistency and narrative coherence, offering a more holistic solution for generating long-form videos.
Multi-Shot Video Generation. Multiple Shot Story Generation focuses on maintaining narrative coherency across distinct scenes, and existing approaches face critical limitations in automation scalability. Animate-a-Story He et al. (2023) uses retrieval-augmented generation to ensure visual consistency but struggles with maintaining logical narrative transitions. MovieDreamer Zhao et al. (2024) requires extensive manual scripting (character profiles, scene details, etc.) and image curation. DreamFactory Xie et al. (2024) demands repetitive adjustment and special documents for each story in its multi-agent system across stories. Flexifilm Ouyang et al. (2024) and StoryDiffusion Zhou et al. (2025) introduce conditional adaptability but still necessitate manual consistency fixes between shots. These methods share a common bottleneck: heavy reliance on human intervention for narrative and visual coherence. VGoT breaks this paradigm through systematic automation: 1) Narrative Fragmentation → LLM-driven story decomposition; 2) Visual Inconsistency → Story-derived identity propagation; 3) Transition Artifacts → Boundary-aware latent processing. This structured approach enables director-level storytelling from single prompts with 10× fewer manual interventions than prior works Zhao et al. (2024); Xie et al. (2024).
@@ 3.1 Preliminaries
Diffusion models Ho et al. (2020); Song et al. (2020a, b) are generative models trained to approximate data distributions $p(x)$ through iterative denoising of random noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. The forward process gradually adds noise according to a variance schedule $\beta_{t}$:
@@
Latent diffusion models Rombach et al. (2022) map this process to a compressed space using a VAE Kingma and Welling (2013), enabling conditional generation through cross-attention:
@@
where $Q=W_{Q}\mathcal{E}(x_{t})$, $K=W_{K}\tau_{\theta}(y)$, and $V=W_{V}\tau_{\theta}(y)$ for text embeddings $y$. Video diffusion models He et al. (2022); Blattmann et al. (2023a) process frame sequences $\{z^{f}\}_{f=0}^{F-1}\in\mathbb{R}^{h\times w\times d}$ through temporal-aware denoising networks $\epsilon_{\theta}(z_{t},t,c)$. FIFO-Diffusion Kim et al. (2024) extends this through queue-based latent processing:
@@
where $\Phi$ denotes the DDIM sampler Song et al. (2020a). While effective for frame continuity, this approach struggles with: (1) abrupt shot transitions that disrupt queue coherence, and (2) integration of image-based conditions. Our framework addresses these through reset mechanisms in adjacent latent-space processing.
@@ 3.2 Problem Definition
Given a one-sentence user input $S$ specifying $N$ shots (e.g., "A story of Mary’s life from birth to death"), we aim to generate a multi-shot video $V$ with reasonable storylines and visual consistency from one sentence through minimal manual intervention. The core challenges are:
@@
  • Consistency: Preserving high-level identity Ye et al. (2023); Yuan et al. (2024); Zhou et al. (2025) while allowing trait variations (e.g., aging expressions, contextual relationships) across shots
@@ 4 Method: VideoGen-of-Thought
In this section, we introduce a structured, step-by-step framework for generating multi-shot videos with reasonable storylines and visual consistency from one sentence with minimal manual intervention, addressing the core challenges of narrative fragmentation, visual inconsistency, and transition artifacts through four distinct yet collaborative modules (Fig 2):
@@ 4.1 Dynamic Storyline Modeling with Self-Vaildation
where $\mathcal{C}_{\text{film}}(p_{\text{cha}},p_{b},p_{r},p_{\text{cam}},p_{h})$ encodes five cinematic dimensions: $p_{\text{cha}}$ governs character appearance evolution and role relationships, $p_{b}$ ensures background consistency across scene transitions, $p_{r}$ maintains interaction patterns and event causality, $p_{\text{cam}}$ specifies shot composition through camera movements, and $p_{h}$ regulates HDR lighting continuity. The self-validation mechanism employs two criteria:
@@
where $C:\mathcal{P}\times\mathcal{P}\rightarrow[0,1]$ measures narrative coherence through pretrained textual feature extractor $E_{\text{GLM}}$ GLM et al. (2024) to compute semantic similarity between consecutive shots, and $K:\mathcal{P}\times\mathcal{C}\rightarrow\{0,1\}$ verifies constraint completeness via rule-based checks against $\mathcal{C}_{\text{film}}$. Thresholds $\tau_{c}=0.85$ and $\tau_{k}=1$ ensure strict adherence to cinematic principles.
@@ 4.2 Identity-Aware Cross-Shot Propagation
where $\mathcal{M}_{I}$ is a pre-trained text-to-image model and $E_{\text{CLIP}}$ denotes CLIP Radford et al. (2021)’s vision encoder. We inject identity features into diffusion via cross-attention:
@@
where $z_{t}\in\mathbb{R}^{n\times d}$ is the latent in step $t$, $\lambda$ balances the influence of text / identity, and $[;]$ denotes concatenation. Keyframe generation integrates both modalities:
@@ 4.3 Adjacent Latent Transition Mechanisms
where $e^{T}_{i}=E_{\text{GLM}}(s_{i})$ uses simplified shot description $s_{i}$ rather than detailed script $p_{i}$, $e^{I}_{i}$ is the keyframe embedding, and $\epsilon_{i}\sim\mathcal{N}(0,\mathbf{I})$ is the initial noise. Inspired by FIFO Kim et al. (2024), we implement boundary-aware noise reset for cross-shot transitions:
@@
EasyAnimate Xu et al. (2024) 0.2402 0.4705 0.0268 0.7969 0.2037
CogVideo Yang et al. (2024) 0.2477 0.6099 0.0222 0.7424 0.2069
VideoCrafter1 Chen et al. (2023) 0.2478 0.3706 0.0350 0.7623 0.1867
VideoCrafter2 Chen et al. (2024) 0.2529 0.5569 0.0686 0.7981 0.1798
@@
EasyAnimate Xu et al. (2024) 0.3333 0.3232 0.3434 0.3535 0.3535 0.3131 0.4646 0.2727 0.2828
CogVideo Yang et al. (2024) 0.1341 0.4146 0.4512 0.2927 0.5976 0.2317 0.1463 0.4512 0.5244
VideoCrafter1 Chen et al. (2023) 0.5446 0.2574 0.1980 0.6436 0.1881 0.1683 0.6535 0.1782 0.1683
VideoCrafter2 Chen et al. (2024) 0.1262 0.4854 0.3883 0.3495 0.3786 0.2718 0.1748 0.4951 0.3981